OpenStudy (anonymous):

help!!!!!!!!!!!!!! Megan exercises every morning for 40 minutes. She does a combination of step aerobics, which burns about 11 calories/min and stretching which burns about 4 calories/min. Her goal is to burn 335 calories during her routine. How long should she participate in each activity in order to burn 335 calories? a. 20 min aerobics 20 min stretching c. 22 min aerobics 18 min stretching b. 30 min aerobics 10 min stretching d. 25 min aerobics 15 min stretching

OpenStudy (anonymous):

a=aerobics s=stretching 335=11a+4s -11a -11a 335-11a=4s divide both side by 4 83.75-2.75a=s 335=11a+83.75-2.75a 335=8.25a+83.75 -83.75 -83.75 251.25=8.25a divide both sides by 8.25 30.45454545454545=a

OpenStudy (anonymous):

its d